Oxford Materials IT Committee Meeting

HILARY TERM 20134
A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 4th February 2014 in the PR Meeting Room.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters oustanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 13th May 2014 at 11am in PR meeting room

ITEM 2. MATTERS OUTSTANDING FROM THE PREVIOUS MEETING

See minutes below.

ITEM 3: MATTERS FOR DISCUSSION

Display Screen Assessment - policy on IT equipment provision
A new round of online DSE assessments is started in January 2014. The DSE assessment promotes the use of height adjustable monitors and separate keyboards and mice in order to achieve correct posture for prolonged computer usage. This means laptop users are being recommended to have external monitor/keyboard/mouse. It is proposed to incorporate this advice into departmental policy that when users purchase laptops as their primary computer they are also advised to purchase external monior/keyboard/mouse.

Verbal reports on the status of various projects:
IT services projects: IT strategy, R12, SITS/BIZTALK, WDE, TONE, maillist upgrade, ICP, ISP Departmental projects: RexRichards, EM, Phones, Photocopiers, iLab

ITEM 4: HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Alex is still providing 40% FTE remote support for the server infrastructure.

Admin Network The migration to Windows7 scheduled for last Christmas is almost ready to be implemented.

Teaching Laboratory The computer room desktops have been (are being) migrated to Windows7. The number of desktops has been reduced from 17 to 12 since the student cohort is now too large to be accommodated in two groups and thus will be split into three groups.

Lecture theatres / meeting rooms All rooms checked at end of term and before start of term. Dawn Hazel is providing weekly board cleaning during term. Failed power supplies on two visualizers were replaced. The Parks Road meeting room AV equipment is to be upgraded for a modest cost.

Digital Print-room Numerous hardware problems with the colour printer have been resolved.

Library The book room computer has failed and will be replaced during the migration to Windows7.

MML The old modelling course desktop computers were all reinstalled as standalone Ubuntu behind a private NAT firewall connecting into a central powerful server for all calculations. Clusters Zener and Landau are to be decommissioned, combined and reinstalled. Rex Richards server room aircon failed again.

EM New private network with storage array has been implemented for the microscopes in Holder Building basement. A similar system is being specified for Begbroke. A Raspberry PI remote desktop solution was implemented for the new Jeol2100. Two new open access desktops have been installed to replace emgroup2&3.

MFFP Due to ceiling collapse numerous computers were relocated. The planned reorganising of the HPC systems is on hold until the ceiling has been repaired.

OMCSimplications of XP end-of-life upon all equipment control systems is still being reviewed.

Website

www.materials and www-solar.materials still need upgrading to the current CMS made simple. Recently live websites: http://nanoanalysis.materials.ox.ac.uk Other sites still under development: none

Publications 5 sets of business cards and several visa passport photographs. New mugshot display boards are being installed to cope with the growth in number of members of department.

Network The network has been suffering intermittent problems which has been affecting the whole department. So far the cause of the problems has not been identified… Two access points have been replaced as the old hardware was no-longer supported by the university management system.

Security No significant infections or security breaches detected.

Disposal Gradual accumulation of old equipment.

Software No significant changes.

Disaster Recovery and Backup The central file servers crashed on XX and 10th Jan but generally uptime has been improved.

Documentation updates Documentation on the private Mediawiki continues to expand.

New Device registrations Nov 23; Dec 7; Jan 23;


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 7th May 2013 in the PR Meeting Room. The following people attended: PJW, RS, KSc, CA, RT, JRY, AL
Apologies: APD, KSu, PE, (SCB)

1. Apologies for Absence and Membership of the Committee

Kumar Sundaram and Phil Edmondson sent their apologies – both away doing experiments. Alana Davies sent her apologies due to family bereavement.

Committee membership was reviewed. Jing Hu is taking over from Andy London as chairman of JCCG.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *Podcasting (Minute 2b)
Podcasting recordings are unlikely to ever happen so this item has been removed.

b) *PC maintenance – OUCS computer hardware breakdown service (Minute 2c)
A process is still required for (automatic?) notification of owners that their warranty is due to expire and directing them to the Equinox hardware breakdown scheme if warranty is required – such as for laptops.

Action: PJW

c) *IT strategic plan (Minute 2d)
The idea of developing a formal strategic plan for IT within the Department of Materials has been shelved for now. We should revisit this idea afresh in 2015.

d) Server Infrastructure upgrade (Minute 2e)
The old fileserver (materials0) is now ready for turn off.

e) *Photocopying and Printing (Minute 2f)
Most networked printers have been created in the new Print server and the drivers tested. Printers will be removed from the old print server when instructions for installing printers have been updated for the new print server. Users are to be notified via the Newsletter and via email about a changeover day which will require all users to reinstall their iPrint printers.

Action: PJW

f) Draft Information Security Policy (Minute 2g)
Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy before Trinity Term.

Action: PJW

The university now has an online information security training module available at http://help.it.ox.ac.uk/infosec/awareness/index which provides guidance to staff about best practice information handling. Member of this IT committee were requested to complete this awareness module and feedback any comments to PJW. In addition PJW will generate a short questionnaire for staff to identify the key information assets and circulate to the ITC for comments.

Action: All/PJW

g) *Finances (Minute 2h)
PJW to circulate financial breakdown for 2010-2011 (and 2011-2012) to the committee.

Action: PJW

h) *EM bookings (Minute 2i)
EG reported that JCCG had requested the equipment booking system be modified to incorporate an instrument status traffic light scheme indicating whether the instrument was fully functional or not. PJW reported that a replacement booking system (iLab) is being selected therefore development of the existing system by PJW has ceased. Andy London asked what the timescale for implementing iLab is and asked whether the existing booking system could continue to be developed in the meantime. PJW will liaise with AL over this request.

Action: PJW

Andy London also asked about whether it would be possible to have per equipment maillists in order to reduce the amount of untargeted emails about microscope sessions becoming available. PJW noted that could be a significant management overhead in doing so. PJW suggested passing this request to Neil Young to see whether the new iLab system has a mechanism for managing user lists.

i) Teaching laboratory computer upgrades (Minute 2j)
The computer room desktops are being migrated to Windows7. The number of desktops has been reduced from 17 to 12 since the student cohort is now too large to be accommodated in two groups and thus will be split into three groups.

j) XP end of life planning (Minute 3a)
Progress has slowed and there is still a lot to do before April 2014.

APD full requested a list of computers still using Windows XP, in particular equipment computers, so that the risks can be reviewed and so that at the next meeting ITC can define a policy concerning what to do with XP computers that cannot be upgraded for whatever reason.

Action: PJW

k) Office 2003 end of life planning (Minute 3b)
The number of computers running Office 2003 (in the past 2 months) remains around 48, all of which are running Windows XP and are thus scheduled to be dealt with as above. IT staff are currently deploying Office2013 on new installations or Office 2010 if required.

RT requested that users upgrading Office should be directed to training courses provided by central IT services and links to online resources to aid the transition. AL noted that Office2013 had been performing well for him.

Action: PJW

l) Novell managed desktops (Minute 3c)
RT queried our current dependency upon Novell. PJW noted that we have used Novell for provision of desktop computing to the teaching laboratory and to administrative staff since the mid 1990s, initially managed via NSMS and subsequently managed by Alex Cock. However the departure of Alex and arrival of Rob who has less experience with Novell provides an opportunity to plan to migrate away from Novell managed desktops. Rob is currently developing an alternative managed desktop environment based upon Microsoft technology. However since the timescale for being able to transition is still unclear the Zenworks upgrade necessary to support Novell with Windows7 is also being implemented by Alex. RT noted the approaching April deadline and was keen to progress with the migration to Windows 7. PJW noted that we are now rolling the teaching lab to Windows7 plus Novell and will start rolling research computers later this month.

Action: RS/AMC

3 Matters for Discussion

a) Display Screen Assessment - policy on IT equipment provision
A new round of online DSE assessments is started in January 2014. The DSE assessment promotes the use of height adjustable monitors and separate keyboards and mice in order to achieve correct posture for prolonged computer usage. This means laptop users are being recommended to have external monitor/keyboard/mouse.

JRY noted that RSI develops over the longer term and typically starts to become an issue when students are in the final stages of thesis writing but the working position need to be addressed much earlier soon after people arrive.

There was discussion about monitors and working positions and the possible benefits of laptop stands as an alternative to an external monitor.

There was agreement that PJW should amend his laptop purchasing advice to recommend also purchasing an external keyboard and mouse and either an external monitor or a laptop stand.

Action: PJW

b) Fresher survey 2013
PJW noted that the results of the annual IT freshers survey have been published. The key points for the department are that students view wifi coverage as essential and top priority. The department should review (and document) current wireless coverage and generate a costed plan for increased coverage, also investigating possible alternative cheaper solutions for providing wireless (currently £600/AP).

Action: KS/RS

4. Head of IT Report
A written report was circulated with the agenda.

5. Finances
No information was presented since R12 transaction data was not available in time for the meeting.

6. AOB

a)EM wireless
AL noted that the wireless coverage in the EM suite is poor. KS reported that the wireless access point had been removed at the request of EM to avoid signal interference with the new microscopes however IT staff challenge the technical justification for this and recommend students to lobby for reconsideration.

7. Date of Next Meeting
The next meeting is on Tuesday 13th May 2014 at 11am in PR meeting room.
The meeting closed at 12.15pm.